OpenStudy (anonymous):

cos(x/5) sin(x/5)=½[sin(2x/5)] true or false?

OpenStudy (lgbasallote):

are you familiar with the identity sin(2a) = 2sinacosa?

OpenStudy (lgbasallote):

let's make it prettier sin(2a) = 2sin(a)cos(a)

OpenStudy (anonymous):

Im not familiar with it.

OpenStudy (lgbasallote):

well your teacher should teach that to you before giving these exercises lol =)) anyway keep that in mind coz im going to use that here and it is used a lot in later maths sin(2a) = 2sin(a)cos(a) now if i replace a by x/5 i will get sin (2x/5) = 2sin(x/5)cos(x/5) agree?

OpenStudy (anonymous):

okay yeah

OpenStudy (lgbasallote):

now if i divide both sides by 2 i will get \[\large \frac{\sin (\frac{2x}{5})}{2} = \sin (\frac{x}{5}) \cos (\frac{x}{5})\] ^the 2 in the right side got canceled

OpenStudy (lgbasallote):

now what can you conclude?

OpenStudy (anonymous):

That its false?

OpenStudy (lgbasallote):

let me rewrite that...i can express this as \[\frac{1}{2} \sin (\frac{2x}{5}) = \sin (\frac{x}{5}) \cos (\frac{x}{5})\]

OpenStudy (lgbasallote):

isn't \[\frac{a}{2} = \frac{1}{2} a?\]

OpenStudy (lgbasallote):

now again..is the equation in your question true? or false?

OpenStudy (anonymous):

The last one makes sense. True.

OpenStudy (lgbasallote):

there you go :D

OpenStudy (anonymous):

thanks (: